OpenStudy (anonymous):

The following is a geometric sequence. 1, -2, 4, -8, 16,…

hartnn (hartnn):

n'th term or sum ?

OpenStudy (anonymous):

True or false

OpenStudy (harsha19111999):

True

hartnn (hartnn):

lol ok find the ratio between each consecutive terms if that ratio is same always, then it is a geometric sequence

hartnn (hartnn):

is -2/1 = 4/-2 = -8/4 = 16/-8 ?? are all those ratios equal ?

hartnn (hartnn):

so you understood, why it is a geometric sequence ?